# Quantum computing with atomic qubit arrays: confronting the cost of connectivity

### v1

@@
Abstract
These notes present a review of the status of quantum computing with arrays of neutral atom qubits, an approach which has demonstrated remarkable progress in the last few years. Scaling digital quantum computing to qubit counts and control fidelities that will enable solving outstanding scientific questions, and provide commercial value, is an outstanding challenge, not least because of the requirement of connecting and entangling distant qubits. long-range Rydberg gates and physical motion outfit atomic qubit arrays with tools for establishing connectivity. These tools operate on different timescales and with distinct levels of parallelization. We analyze several prototypical architectures from the perspective of achieving fast connectivity for circuits with large scale entanglement, as well as fast cycle times for measurement based quantum error correcting codes. Extending Rydberg interactions to multiple atomic species has emerged as a promising route to achieving this latter requirement.
I Introduction
@@
The concept of a quantum computer is not new[1, 2, 3, 4], yet the timeline for wide availability of quantum computers that are useful for solving problems in science or commerce is uncertain.111 The material presented here is based on lectures given at the International School of Physics “Enrico Fermi” in Varenna, Italy in July, 2024. Given the availability of other review papers and the very rapid pace of development of quantum computing with neutral atoms I have chosen not to include all the background material presented in the lectures. Instead some new developments up until Spring 2025 when these notes were finalized are presented. The uncertainty does not arise from questions about the predictive power of quantum mechanics, which was established 100 years ago and has been verified in countless experiments. Rather, scientists and engineers who are striving to build quantum computers do not yet have a verified blueprint for how to proceed. It should also be acknowledged that objections to the completeness of quantum mechanics[5], and its seemingly absurd predictions about the physical world[6, 7, 8] have persisted[9], despite the success of experiments which have confirmed the violation of Bell’s inequality[10, 11, 12] as predicted by quantum theory[13]. One of the fundamental applications of large scale quantum computing may well be the creation of new tools for studying the foundations of quantum theory.
@@
There is no single reason why developing a quantum computer is so challenging. Even so, a few obstacles stand out. Quantum computational advantage derives from the possibility of controlling quantum phenomena including superposition and entanglement. Preserving these quantum phenomena demands that the qubits are well isolated from the environment. On the other hand a computer is not just a science experiment but is a machine that can be programmed to perform desired tasks, which requires control. Achieving both isolation and control in a single system is difficult and has been tackled most successfully in just a handful of physical platforms, including trapped ions, superconducting circuits, photonics, electrons in low-dimensional semiconductors, and neutral atoms[14]. Apart from the dichotomy of isolation and control there is the challenge of connectivity. In conventional computers information is freely copied and moved across circuits on wires. In the quantum world unknown states cannot be copied[15, 16], and can only be moved to a different physical location if no trace of the state is left behind. As a consequence performing logic with qubits that may be physically far apart is not trivial and significantly adds to the challenge of scaling up quantum computers. This challenge is present in all of the aforementioned platforms. Neutral atoms provide two mechanisms for achieving connectivity: long-range Rydberg interactions[17, 18], and physical transport of qubits[19]. These two mechanisms operate on different spatial and temporal scales which has consequences for how to optimize the connectivity architecture.
@@
Before proceeding to specifics of neutral atom qubits, it seems useful to delineate two distinct paths forward, and the implications for future progress. These paths are indicated in Fig. 1. The current state of play has been termed the NISQ era for Noisy Intermediate Scale Quantum[20]. Working quantum computers exist and have demonstrated impressive performance, even though they have a limited number of qubits and suffer from errors in control and measurements. Despite these limitations NISQ machines have performed calculations based on random circuit sampling (RCS) that produce bitstrings with a statistical distribution that cannot be mimicked by conventional computers. These so-called “Quantum Supremacy" experiments[21, 22, 23, 24, 25] demonstrate the capability of performing a particular calculation that is beyond the reach of conventional computers. but not yet a calculation that has wider utility. The possibility of using RCS not just to demonstrate the power of a quantum computer, but to solve useful problems, has not been decisively demonstrated, although there are promising opportunities involving applications of certified randomness[26] and quantum fingerprinting for comparing remote databases[27]. Other demonstrations of NISQ hardware that seek to demonstrate computational advantage for digital quantum circuits have targeted the transverse field Ising model[28, 29] which has broad applicability as a description of interactions and dynamics of many-body spin systems. Very recent results enabled by high fidelity entangling gates simulated prethermalization in a regime beyond currently available classical techniques[29].
While the progress in RCS and transverse field Ising simulations with quantum processors is remarkable, the progress has also inspired development of classical algorithms that successfully competed with early quantum demonstrations[30, 31]. This situation with a back and forth competition between conventional computers and digital quantum hardware begs the question of the viability of quantum computational advantage in the current NISQ era without quantum error correction and fault tolerance.
To clarify this question we need to establish a connection between the fidelity of expectation values of observables provided by a quantum circuit and the classical cost of simulating the observable. The performance of state of the art NISQ processors is typically limited by the fidelity of the two-qubit gate used for entanglement. As discussed in [32] the expectation value of an operator $\hat{O}$ that is obtained from a noisy quantum state with density matrix $\rho$ can be expressed as ${\rm tr}(\rho\hat{O})={\mathcal{F}}_{\rm eff}\langle\hat{O}\rangle_{\rm ideal}$ where ${\mathcal{F}}_{\rm eff}$ is the effective fidelity of the observable and $\langle\hat{O}\rangle_{\rm ideal}$ is the noise free expectation value. An estimate for the fidelity is
@@
where $\epsilon$ is the two-qubit gate infidelity and $V_{\rm eff}$ is the circuit volume, which may be approximated by the number of two-qubit gates involved in calculating the expectation value. A tensor network simulation[33] of the expectation value on a classical computer has a computational cost that approximately scales as
@@
The scaling of $A$ with the number of qubits and gate fidelity depends on a number of factors including the dimensionality and connectivity of the circuit, the type of entangling gate used, and the physical problem of interest. For RCS and short evolution times $A\sim\sqrt{n}t$ with $t$ the execution time (circuit depth) for a square array of $n$ qubits[32]. The scaling relations (1,2) imply that a noisy circuit will output a fidelity that decreases exponentially with circuit volume, while the classical cost of simulating the result increases exponentially with the area of a circuit cut, which can be thought of as a proxy for the maximum amount of entanglement generated. Crucially the exponential scaling of the fidelity and the classical computational cost are different[32]. Putting $t\sim 1/\epsilon$ we have $C\sim 2^{\sqrt{n}/\epsilon}$ which has been visualized in Fig. 2. Experimental progress in the last few years has led to a rapid increase of the estimated equivalent computational cost of RCS performed on quantum hardware. For that reason recent papers[23, 25] choose to present $\log(\log(C))$ and we have made the same choice in Fig. 2. It can be seen that the contours of constant $C$ make a small angle with the abscissa and have a large gradient aligned towards the ordinate. Thus the critical requirement for increasing the computational cost is the gate fidelity, not the number of qubits. This is consistent with other measures of computational power such as the quantum volume[34].
@@
State of the art utility experiments are being performed with $n\sim 100$ and $\epsilon=1-{\mathcal{F}}\sim 0.001.$ It is apparent from Fig. 2 that increasing the qubit count up to 1000 has only a moderate impact on the computational cost. On the other hand improving the gate fidelity by another factor of 10 will give a much more pronounced boost to the cost of a classical simulation. The implication is that the path to HP-NISQ suggested in Fig. 1 primarily requires increasing control fidelity. Realizing the full potential of quantum computers will undoubtedly require FTQC with orders of magnitude more qubits than HP-NISQ machines. While higher gate fidelity will help to reduce the requirements on qubit count, even at extremely high fidelities the code rate will at best be $1/5$[35] with fault tolerant universality needing additional qubits for magic state distillation[36].
@@
The rest of this contribution is organized as follows. In Sec. II we give a brief summary of the current state of the art for neutral atom, gate based quantum computers. Progress in scaling qubit count, and performing fast, high fidelity measurements is reviewed. As described above higher gate fidelity is crucial for realizing computational power without error correction. Some fundamental limits and prospects for increasing neutral atom gate fidelity are discussed in Sec. III. Operation at the logical level with many more physical qubits brings the challenge of connectivity to the forefront. Several approaches are compared in Sec. IV. An additional requirement for logical qubit operations is mid-circuit measurements which are discussed with a focus on a two-species architecture in Sec. V. We conclude with a perspective on the present challenges and an outlook for future developments in Sec. VI.
II Status and scaling of neutral atom qubit arrays
@@
In the last few years there has been a remarkable and rapid increase in the rate of development of neutral atom quantum processors. Many new academic research groups have entered the field and there are more than five companies seeking to commercialize neutral atom quantum computing. Key performance metrics of state of the art systems include large 2D arrays with hundreds of qubits, qubit coherence times of several seconds, gate operation times of a few $\mu\rm s$ to sub-$\mu\rm s$ for one- and two-qubit gates, single shot measurement and state preparation and measurement (SPAM) fidelity above 99%. Gate fidelities have reached ${\mathcal{F}}>0.9999$ for global rotations driven by microwaves[37, 38], ${\mathcal{F}}>0.999$ for qubit specific single qubit gates[39, 40], and ${\mathcal{F}}>0.99$ for two-qubit gates. Using long-range Rydberg interactions or atom motion, connectivity can extend beyond nearest neighbor as will be discussed in detail in Sec. IV.
@@
With these advances neutral atom qubit arrays are now positioned as a leading modality in the race to realize useful quantum computing. Arrays with hundreds of qubits have been used for analog simulation of interacting spin models and solution of optimization problems[41, 42, 43]. Gate model quantum circuits have demonstrated basic algorithms at the physical qubit level[44, 40] as well as circuits with logical qubits[19, 45, 46, 47], circuit based preparation of entangled states for metrology[48, 49], and first demonstrations of magic state distillation[50]. Capabilities that are being actively improved upon include qubit counts, fast and low loss qubit measurements, and gate fidelities.
@@
In terms of qubit counts there have been several demonstrations of atom arrays with more than 1000 trapping sites[42, 51, 52, 53, 54] with the latest results achieving 2D arrays with 6100 atoms loaded on average into 12000 sites[55], and 2000 atoms deterministically rearranged into desired patterns with almost no defects[56]. The Talbot effect[57] leads to axial repetition of transversally periodic structures and can be an annoyance when preparing a single layer of atoms, but can also be advantageous for preparing multiple, well spaced trapping planes without requiring additional laser power[58]. Atom array assembly into three dimensional structures is also possible[59, 60, 56, 61]. In principle extremely large arrays could be prepared in a single vacuum chamber. Depending on the atom used, the wavelength of the trapping light, and the desired trap depth as little as 1 mW of optical power is required for each site. Thus $10^{6}$ sites would require 1 kW of optical power. At an atom spacing of $5~{}\mu\rm m$ the array would cover an area of only 5 mm on a side. Lasers of this power level are available, although integration with high resolution lens assemblies presents challenges related to thermal management, and mitigation of background light scattering.
@@
Figure 3: Measurement of neutral atom qubit states in free space. The results are quantified in terms of the measurement time and the generalized fidelity given by the state detection fidelity $F$ times the atom retention probabiity $1-P_{\rm loss}$ The chronological sequence of these results follows the alphabetical ordering. The experiments a,b,g used single photon counting detectors and all others used cameras: a[62], b[63], c[64], d[65], e[66], f[67], g[68], h[53], i[69], j[70], k[71].
@@
An important consideration when preparing very large arrays without defects using atom rearrangement[72, 73, 74] is the requirement of measuring whether or not an atom is present while retaining the atom for subsequent use after the measurement. In order to prepare an array of $N$ atoms with no defects occupancy measurements with an accuracy and loss less than $1/N$ are required. Atom occupancy measurements with a fidelity of $F=0.99991(1)$ and a loss probability of $P_{\rm loss}=1-0.99932(8)$ were demonstrated in [67]. These results support preparation of fully occupied arrays with thousands of atoms provided the array assembly time is short compared to the vacuum limited lifetime.
@@
Within quantum circuits it is necessary to measure not just the presence of an atom but the quantum state. Figure 3 presents qubit measurements performed by imaging state dependent fluorescence light. It is notable that the results with the best combination of fidelity and atom retention also required the longest integration times. This is because imaging faster requires a higher scattering rate which leads to more heating and atom loss. High fidelity and very fast measurements are possible using optical cavities[75, 76, 77, 78], but integration of cavities with large atom arrays remains a challenge.
III Rydberg gate fidelity

### v3

@@
11institutetext: Department of Physics, University of Wisconsin-Madison, Madison, WI 53706, USA 22institutetext: Infleqtion, Madison, WI, 53703 USA, 22email: msaffman@wisc.edu
Quantum computing with atomic qubit arrays: confronting the cost of connectivity
[…]
1 Introduction
@@
The concept of a quantum computer is not new Feynman1982 ; Feynman1986 ; Manin1980 ; Deutsch1985 , yet the timeline for wide availability of quantum computers that are useful for solving problems in science or commerce is uncertain. [^1] The material presented here is based on lectures given at the International School of Physics “Enrico Fermi” in Varenna, Italy in July, 2024. Given the availability of other review papers and the very rapid pace of development of quantum computing with neutral atoms I have chosen not to include all the background material presented in the lectures. Instead some new developments up until summer 2025 when these notes were finalized are presented. The uncertainty does not arise from questions about the predictive power of quantum mechanics, which was established 100 years ago and has been verified in countless experiments. Rather, scientists and engineers who are striving to build quantum computers do not yet have a verified blueprint for how to proceed. It should also be acknowledged that objections to the completeness of quantum mechanics Einstein1935 , and its seemingly absurd predictions about the physical world Schrodinger1935 ; Schrodinger1935b ; Schrodinger1936 have persisted tHooft2021 , despite the success of experiments which have confirmed the violation of Bell’s inequality Giustina2015 ; Shalm2015 ; Rosenfeld2017 as predicted by quantum theory Bell1964 . One of the fundamental applications of large scale quantum computing may well be the creation of new tools for studying the foundations of quantum theory.
@@
There is no single reason why developing a quantum computer is so challenging. Even so, a few obstacles stand out. Quantum computational advantage derives from the possibility of controlling quantum phenomena including superposition and entanglement. Preserving these quantum phenomena demands that the qubits are well isolated from the environment. On the other hand a computer is not just a science experiment but is a machine that can be programmed to perform desired tasks, which requires control. Achieving both isolation and control in a single system is difficult and has been tackled most successfully in just a handful of physical platforms, including trapped ions, superconducting circuits, photonics, electrons in low-dimensional semiconductors, and neutral atoms Bergou2021 . Apart from the dichotomy of isolation and control there is the challenge of connectivity. In conventional computers information is freely copied and moved across circuits on wires. In the quantum world unknown states cannot be copied Wootters1982 ; Dieks1982 , and can only be moved to a different physical location if no trace of the state is left behind. As a consequence performing logic with qubits that may be physically far apart is not trivial and significantly adds to the challenge of scaling up quantum computers. This challenge is present in all of the aforementioned platforms. Neutral atoms provide two mechanisms for achieving connectivity: long-range Rydberg interactions Lukin2001 ; Saffman2010 , and physical transport of qubits Bluvstein2022 . These two mechanisms operate on different spatial and temporal scales which has consequences for how to optimize the connectivity architecture.
@@
Before proceeding to specifics of neutral atom qubits, it seems useful to delineate two distinct paths forward, and the implications for future progress. These paths are indicated in Fig. 1. The current state of play has been termed the NISQ era for Noisy Intermediate Scale Quantum Preskill2018 . Working quantum computers exist and have demonstrated impressive performance, even though they have a limited number of qubits and suffer from errors in control and measurements. Despite these limitations NISQ machines have performed calculations based on random circuit sampling (RCS) that produce bitstrings with a statistical distribution that cannot be mimicked by conventional computers. These so-called “Quantum Supremacy" experiments Arute2019 ; YWU2021 ; Morvan2024 ; DeCross2025 ; DGao2025 demonstrate the capability of performing a particular calculation that is beyond the reach of conventional computers. but not yet a calculation that has wider utility. The possibility of using RCS not just to demonstrate the power of a quantum computer, but to solve useful problems, has not been decisively demonstrated, although there are promising opportunities involving applications of certified randomness MLiu2025 and quantum fingerprinting for comparing remote databases Gokhale2022 . Other demonstrations of NISQ hardware that seek to demonstrate computational advantage for digital quantum circuits have targeted the transverse field Ising model YKim2023 ; Haghshenas2025 which has broad applicability as a description of interactions and dynamics of many-body spin systems. Very recent results enabled by high fidelity entangling gates simulated prethermalization in a regime beyond currently available classical techniques Haghshenas2025 .
While the progress in RCS and transverse field Ising simulations with quantum processors is remarkable, the progress has also inspired development of classical algorithms that successfully competed with early quantum demonstrations FPan2022 ; Begusic2024 . This situation with a back and forth competition between conventional computers and digital quantum hardware begs the question of the viability of quantum computational advantage in the current NISQ era without quantum error correction and fault tolerance.
To clarify this question we need to establish a connection between the fidelity of expectation values of observables provided by a quantum circuit and the classical cost of simulating the observable. The performance of state of the art NISQ processors is typically limited by the fidelity of the two-qubit gate used for entanglement. As discussed in Kechedzhi2024 the expectation value of an operator $\hat{O}$ that is obtained from a noisy quantum state with density matrix $\rho$ can be expressed as ${\rm tr}(\rho\hat{O})={\mathcal{F}}_{\rm eff}\langle\hat{O}\rangle_{\rm ideal}$ where ${\mathcal{F}}_{\rm eff}$ is the effective fidelity of the observable and $\langle\hat{O}\rangle_{\rm ideal}$ is the noise free expectation value. An estimate for the fidelity is
@@
where $\epsilon$ is the two-qubit gate infidelity and $V_{\rm eff}$ is the circuit volume, which may be approximated by the number of two-qubit gates involved in calculating the expectation value. A tensor network simulation Orus2014 of the expectation value on a classical computer has a computational cost that approximately scales as
@@
The scaling of $A$ with the number of qubits and gate fidelity depends on a number of factors including the dimensionality and connectivity of the circuit, the type of entangling gate used, and the physical problem of interest. For RCS and short evolution times $A\sim\sqrt{n}t$ with $t$ the execution time (circuit depth) for a square array of $n$ qubits Kechedzhi2024 . The scaling relations (1,2) imply that a noisy circuit will output a fidelity that decreases exponentially with circuit volume, while the classical cost of simulating the result increases exponentially with the area of a circuit cut, which can be thought of as a proxy for the maximum amount of entanglement generated. Crucially the exponential scaling of the fidelity and the classical computational cost are different Kechedzhi2024 . Putting $t\sim 1/\epsilon$ we have $C\sim 2^{\sqrt{n}/\epsilon}$ which has been visualized in Fig. 2. Experimental progress in the last few years has led to a rapid increase of the estimated equivalent computational cost of RCS performed on quantum hardware. For that reason recent papers Morvan2024 ; DGao2025 choose to present $\log(\log(C))$ and we have made the same choice in Fig. 2. It can be seen that the contours of constant $C$ make a small angle with the abscissa and have a large gradient aligned towards the ordinate. Thus the critical requirement for increasing the computational cost is the gate fidelity, not the number of qubits. This is consistent with other measures of computational power such as the quantum volume Cross2019 .
@@
State of the art utility experiments are being performed with $n\sim 100$ and $\epsilon=1-{\mathcal{F}}\sim 0.001.$ It is apparent from Fig. 2 that increasing the qubit count up to 1000 has only a moderate impact on the computational cost. On the other hand improving the gate fidelity by another factor of 10 will give a much more pronounced boost to the cost of a classical simulation. The implication is that the path to HP-NISQ suggested in Fig. 1 primarily requires increasing control fidelity. Realizing the full potential of quantum computers will undoubtedly require FTQC with orders of magnitude more qubits than HP-NISQ machines. While higher gate fidelity will help to reduce the requirements on qubit count, even at extremely high fidelities the code rate will at best be $1/5$ Laflamme1996 with fault tolerant universality needing additional qubits for magic state distillation Bravyi2005 .
@@
The rest of this contribution is organized as follows. In Sec. 2 we give a brief summary of the current state of the art for neutral atom, gate based quantum computers. Progress in scaling qubit count, and performing fast, high fidelity measurements is reviewed. As described above higher gate fidelity is crucial for realizing computational power without error correction. Some fundamental limits and prospects for increasing neutral atom gate fidelity are discussed in Sec. 3. Operation at the logical level with many more physical qubits brings the challenge of connectivity to the forefront. Several approaches are compared in Sec. 4. An additional requirement for logical qubit operations is mid-circuit measurements which are discussed with a focus on a two-species architecture in Sec. 5. We conclude with a perspective on the present challenges and an outlook for future developments in Sec. 6.
2 Status and scaling of neutral atom qubit arrays
@@
In the last few years there has been a remarkable and rapid increase in the rate of development of neutral atom quantum processors. Many new academic research groups have entered the field and there are more than five companies seeking to commercialize neutral atom quantum computing. Key performance metrics of state of the art systems include large 2D arrays with hundreds of qubits, qubit coherence times of several seconds, gate operation times of a few $\mu\rm s$ to sub-$\mu\rm s$ for one- and two-qubit gates, single shot measurement and state preparation and measurement (SPAM) fidelity above 99%. Gate fidelities have reached ${\mathcal{F}}>0.9999$ for global rotations driven by microwaves CSheng2018 ; Nikolov2023 , ${\mathcal{F}}>0.999$ for qubit specific single qubit gates Radnaev2025 ; Chinnarasu2025 , and ${\mathcal{F}}>0.99$ for two-qubit gates. Using long-range Rydberg interactions or atom motion, connectivity can extend beyond nearest neighbor as will be discussed in detail in Sec. 4.
@@
With these advances neutral atom qubit arrays are now positioned as a leading modality in the race to realize useful quantum computing. Arrays with hundreds of qubits have been used for analog simulation of interacting spin models and solution of optimization problems Scholl2021 ; Ebadi2021 ; Ebadi2022 . Gate model quantum circuits have demonstrated basic algorithms at the physical qubit level Graham2022 ; Chinnarasu2025 as well as circuits with logical qubits Bluvstein2022 ; Bluvstein2024 ; Bedalov2024 ; Reichardt2024 , circuit based preparation of entangled states for metrology ACao2024 ; Finkelstein2024 , and first demonstrations of magic state distillation Rodriguez2025 . Capabilities that are being actively improved upon include qubit counts, fast and low loss qubit measurements, and gate fidelities.
@@
In terms of qubit counts there have been several demonstrations of atom arrays with more than 1000 trapping sites Ebadi2021 ; Huft2022 ; Pichard2024 ; Norcia2024 ; Gyger2024 with the latest results achieving 2D arrays with 6100 atoms loaded on average into 12000 sites Manetsch2025 , and 2000 atoms deterministically rearranged into desired patterns with almost no defects RLin2025 . The Talbot effect Talbot1836 leads to axial repetition of transversally periodic structures and can be an annoyance when preparing a single layer of atoms, but can also be advantageous for preparing multiple, well spaced trapping planes without requiring additional laser power Schlosser2023 . Atom array assembly into three dimensional structures is also possible Barredo2018 ; Kumar2018 ; RLin2025 ; Kusano2025 . In principle extremely large arrays could be prepared in a single vacuum chamber. Depending on the atom used, the wavelength of the trapping light, and the desired trap depth as little as 1 mW of optical power is required for each site. Thus $10^{6}$ sites would require 1 kW of optical power. At an atom spacing of $5\penalty 10000\ \mu\rm m$ the array would cover an area of only 5 mm on a side. Lasers of this power level are available, although integration with high resolution lens assemblies presents challenges related to thermal management, and mitigation of background light scattering.
@@
Figure 3: Representative measurements of neutral atom qubit states in free space reported in the literature. The results are quantified in terms of the measurement time and the generalized fidelity given by the state detection fidelity $F$ times the atom retention probability $1-P_{\rm loss}$ . The experiments a,b,g,h with the shortest integration times used single photon counting detectors and all others used cameras: a Fuhrmanek2011 , b Gibbons2011 , c Kwon2017 , d Martinez-Dorantes2017 , e TYWu2019 , f Covey2019a , g Shea2020 , h Chow2023 , i Huie2023 , j Lis2023 , k Norcia2024 , l RTao2024 , m SMa2023 , n Scott2025 .
@@
An important consideration when preparing very large arrays without defects using atom rearrangement WLee2016 ; Barredo2016 ; Endres2016 is the requirement of measuring whether or not an atom is present while retaining the atom for subsequent use after the measurement. In order to prepare an array of $N$ atoms with no defects occupancy measurements with an accuracy and loss less than $1/N$ are required. Atom occupancy measurements with a fidelity of $F=0.99991(1)$ and a loss probability of $P_{\rm loss}=1-0.99932(8)$ were demonstrated in Covey2019a . These results support preparation of fully occupied arrays with thousands of atoms provided the array assembly time is short compared to the vacuum limited lifetime.
@@
Within quantum circuits it is necessary to measure not just the presence of an atom but the quantum state. Figure 3 presents qubit measurements performed by imaging state dependent fluorescence light. It is notable that the results with the best combination of fidelity and atom retention also required the longest integration times. This is because imaging faster requires a higher scattering rate which leads to more heating and atom loss. High fidelity and very fast measurements are possible using optical cavities Bochmann2010 ; Volz2011 ; Deist2022 ; Grinkemeyer2025 , but integration of cavities with large atom arrays remains a challenge.
3 Rydberg gate fidelity